OpenStudy (republican31):

Solve Sqrt(4y+5)-Sqrt(y-1)=3

OpenStudy (republican31):

Please tell me if I'm headed in the right direction and guide me please!

OpenStudy (republican31):

Sqrt(4y+5)-Sqrt(y-1)=3 Sqrt(4y+5)=3+Sqrt(y-1) 4y+5=9+6Sqrt(y-1)+y-1

OpenStudy (republican31):

4y+5=8+6Sqrt(y-1)+y 3y+5=8+6Sqrt(y-1) 3y-3=6Sqrt(y-1)

OpenStudy (republican31):

1/2y-1/2=Sqrt(y-1) 1/4y^2-2/4y-1/4=y-1 y^2-2y-1=y-1

OpenStudy (republican31):

y^2-2y=y y^2-3y=0 y(y-3)=0 y={0,3} but that's not correct, it's {1,5}

TheSmartOne (thesmartone):

you should get \(\Large\frac{1}{4}y^2 - \frac{1}{2}y \color{red}{+}\frac{1}{4}=y-1\)

TheSmartOne (thesmartone):

(a - b)^2 = a^2 - 2ab + b^2 it should be a plus, not a minus also I just went ahead and simplified 2/4y to 1/2y

OpenStudy (republican31):

I would have but I thought I would multiply by 4 to get rid of the fractions. And thanks I didn't catch that mistake.

OpenStudy (republican31):

So revised: 1/4y^2-2/4y+1/4 y^2-2y+1=y-1 y^2-2y+2=y

OpenStudy (republican31):

y^2-3y+2=0 (y-1)(y-2)=0 y={1,2} still not {1,5} so I'm still wrong somewhere.

TheSmartOne (thesmartone):

If a = b then ac = bc you multiplied 4 on the left side, you have to do so at the right side also

OpenStudy (republican31):

Oh....

OpenStudy (republican31):

Thank you so much!!!! Now off to reattempt that test!!!! THANK YOU!

OpenStudy (republican31):

YESSS I GOT IT RIGHT THANK YOU SOOOOO MUCH YOU HAVE NO IDEA HOW MUCH I ALMOST CRIED OVER THIS ONE